OpenStudy (anonymous):

Which has the greatest value? 5 out of 9 zero point five thirty percent three-eighths @mathstudent55

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@melanielovee4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Melodysim

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

what's 5/9?

OpenStudy (anonymous):

5 divided by 9

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

0.5 = 1/2 if you compare 5/9 and 1/2 set them with same denominator 5/9 = 10/18 1/2 = 9/18 which ones is bigger? 10/18 or 9/18

OpenStudy (anonymous):

10/18

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

If it were me, I'd change them all to decimals

OpenStudy (anonymous):

10/18= 5/9 so you know that 5/9 > 0.5 30 percent = 30/100 = 0.3, so it's smaller than 0.5 So far 5/9 > 0.5 > 0.3

OpenStudy (anonymous):

5/9=.55 zero point five=.5 thirty percent= .3 3/8= .37

OpenStudy (anonymous):

so yes like @1lorax2 say 5/19 is biggest. :)) <3

OpenStudy (anonymous):

Which is biggest?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you @study100

OpenStudy (anonymous):

do you know how to turn fractions to decimals though? I assume you can't use calculators

OpenStudy (anonymous):

ohhhh....without calculators, fractions can become difficult. dont you multiply the fraction by 1/100?

OpenStudy (anonymous):

I'm not entirely sure.